Table 1            Estimated Total Returns                                                                                                 
State of the Economy Probability T-Bond SETX Golden S&P 500
Recession 5% 5% -19% 20% -14%
Below Average 15% 5% 2% 13% 3%
Average 45% 5% 9% 10% 11%
Above Average 25% 5% 34% 5% 22%
Boom 10% 5% 25% -5% 33%

SETX expects to have $2.72 in earnings next year. SETX had $1.95 in earnings 4 years ago. What should SETX sell for? If it sells for $25, should you buy?
Answer #1

Growth Rate = (2.72/1.95)1/4 - 1 = 8.97%

Expected Rate = 0.05(-0.19) + 0.15(0.02) + 0.45(0.09) + 0.25(0.34) + 0.10(0.25)

Expected Rate = 14.40%

Using DDM model,

Stock Price = 2.72/(0.1440 - 0.0897)

Stock Price = $50.09

As it sells for $25, one should buy the stock.